OpenStudy (anonymous):

find the slope of the line passing through the given points: (1/2, 3/5) and (1/4,1/3)

OpenStudy (anonymous):

\[\text{slope} = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

\[(x_1, y_1) = \left(\frac{1}{2}, \frac{3}{5}\right)\]\[(x_2, y_2) = \left(\frac{1}{4}, \frac{1}{3}\right)\]

OpenStudy (anonymous):

You should be able to do the rest.